OpenStudy (anonymous):

How to create a quadratic equation when given is that one of the solutions must be 1/(3+2sqrt2)

OpenStudy (mertsj):

If one solution is 1/(3+2sqrt2 then there must be another root of 1/3-2sqrt2 since they always come in pairs

OpenStudy (king):

yeah mertsj is rite..but mertsj when roots are real they are not in pairs

OpenStudy (mertsj):

Remember the quadratic formula \[x=\frac{-b \pm \sqrt{b^2-4ac}}{2a}\]

OpenStudy (mertsj):

See that plus or minus thing? That's why they always come in pairs.

OpenStudy (experimentx):

i think we should rationalize first

OpenStudy (anonymous):

Can you show all the steps?

OpenStudy (mertsj):

@king If roots are irrational they are in pairs.

OpenStudy (king):

yeah exactly what i was sayin!

OpenStudy (mertsj):

\[\frac{1}{3+2\sqrt{2}}\times\frac{3-2\sqrt{2}}{3-2\sqrt{2}}=\frac{3-2\sqrt{2}}{9-8}=3-2\sqrt{2}\]

OpenStudy (anonymous):

Alright, no idea.

OpenStudy (mertsj):

@king It's not what you were saying. You said if the roots are real, they are not in pairs and irrational numbers are real numbers.

OpenStudy (mertsj):

so Erko. One of the roots is 3-2sqrt2. That means the other one is 3+2sqrt2

OpenStudy (king):

@Mertsj sry i meant that only!

OpenStudy (mertsj):

So the equation is x minus the first root times x - the second root

OpenStudy (mertsj):

\[[x-(3-2\sqrt{2})][x-(3+2\sqrt{2})]=0\]

OpenStudy (mertsj):

and there is your equation.